OpenStudy (anonymous):

Question 1 (Multiple Choice Worth 4 points) [9.02] What are the x-intercept(s) of the graph of y − x2 = −11x + 24? (−3, 0) and (8, 0) (3, 0) and (−8, 0) (−3, 0) and (−8, 0) (3, 0) and (8, 0)

OpenStudy (anonymous):

Help me please

jimthompson5910 (jim_thompson5910):

Plug in y = 0 and solve for x \[\large y - x^2 = -11x + 24\] \[\large 0 - x^2 = -11x + 24\] \[\large -x^2 = -11x + 24\] \[\large -x^2+x^2 = -11x + 24+x^2\] \[\large 0 = x^2-11x+24\] \[\large x^2-11x+24 = 0\] I'll let you finish up

OpenStudy (anonymous):

What do i do next i am lost sorry

jimthompson5910 (jim_thompson5910):

factoring is the next step

OpenStudy (anonymous):

I don't know how to do that

jimthompson5910 (jim_thompson5910):

find two numbers that multiply to 24 and add to -11

OpenStudy (anonymous):

can you do this one and i next question is that ok

jimthompson5910 (jim_thompson5910):

what two numbers multiply to 24 and add to -11?

jimthompson5910 (jim_thompson5910):

list out the factors of 24 and see which two add to -11

OpenStudy (anonymous):

I am lost

jimthompson5910 (jim_thompson5910):

list out all the possible ways to multiply to 24

jimthompson5910 (jim_thompson5910):

so you'll have 1*24, 2*12, etc etc

OpenStudy (anonymous):

oh ok. 3x8= 24

OpenStudy (anonymous):

2x12=24

OpenStudy (anonymous):

4x6=24

OpenStudy (anonymous):

6x4=24

jimthompson5910 (jim_thompson5910):

good, notice how 3+8 = 11, 2+12 = 14, 4+6 = 10 none of these results are -11

jimthompson5910 (jim_thompson5910):

so how do we fix that?

OpenStudy (anonymous):

I don't know how

jimthompson5910 (jim_thompson5910):

well notice how -1 times -24 = +24 or just 24 and how -2 times -12 = +24 or just 24 etc etc

OpenStudy (anonymous):

-6x24=-24

OpenStudy (anonymous):

Is this the answers x - 3 = 0 or x - 8 = 0

OpenStudy (anonymous):

or this one x = 3 or x = 8

OpenStudy (anonymous):

is answers B

jimthompson5910 (jim_thompson5910):

good, x^2 - 11x + 24 factors to (x-3)(x-8)

OpenStudy (anonymous):

so its B

OpenStudy (anonymous):

And can you help me on this one Question 2 (Multiple Choice Worth 4 points) [9.02] What are the solutions of x2 − 3x − 28 = 0? x = 4, x = 7 x = −4, x = −7 x = 4, x = −7 x = −4, x = 7

OpenStudy (anonymous):

Is this right x^2-3x-28=0 (x+4)(x-7)=0 x+4=0 x=-4 x-7=0 x=7 So, x could be -4 or 7.

jimthompson5910 (jim_thompson5910):

perfect

OpenStudy (anonymous):

is it this one x = −4, x = 7 or this x = 4, x = −7

jimthompson5910 (jim_thompson5910):

you tell me

OpenStudy (anonymous):

I think its this one x = 4, x = −7

jimthompson5910 (jim_thompson5910):

but you said "So, x could be -4 or 7."

OpenStudy (anonymous):

I don't know one of them

jimthompson5910 (jim_thompson5910):

one of them has x=4, the other has x =-4 so clearly they are different

OpenStudy (anonymous):

yea

jimthompson5910 (jim_thompson5910):

so I'm sure you see which one is the answer